OpenStudy (anonymous):

Please help, will medal 2a-9 4a+4 —— • —— 2a-6 a+3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

see if you can factor and cancel some stuff

OpenStudy (misty1212):

for example, \(4a+4=4(a+1)\)

OpenStudy (anonymous):

oh ok just a sec

OpenStudy (misty1212):

how about \(2a-6\) can you factor that one?

OpenStudy (anonymous):

ya ok so it would be 2a-6 = 2(a-3) ?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

but not much cancels here anways, just a little

OpenStudy (anonymous):

so we've got 2a-9 4(a+1) —— • ——- 2(a-3) a+3

OpenStudy (misty1212):

you get \[\frac{2a-9)4(x+1)}{2(x-3)(x+3)}\]

OpenStudy (misty1212):

ooh i got the questions marks of death, let me refresh and see if they go away

OpenStudy (anonymous):

whoa ya that

OpenStudy (misty1212):

kk so all that cancels is a lousy 2 `

OpenStudy (misty1212):

\[\frac{(2a-9)4(x+1)}{2(x-3)(x+3)}=\frac{(2a-9)2(x+1)}{(x-3)(x+3)}\]

OpenStudy (misty1212):

i guess if you have to you must multiply this mess out

OpenStudy (anonymous):

I believe that is what they want...

OpenStudy (misty1212):

then if you have to , you have to the denominator is easy right?

OpenStudy (misty1212):

if not let me know and i can show you what it is

OpenStudy (anonymous):

I'm sorry, I do really poorly at this :( please explain...

OpenStudy (misty1212):

ok lets do the bottom first

OpenStudy (misty1212):

\[(a+3)(a-3)=a^2+3a-3a+9=a^2-9\]

OpenStudy (misty1212):

or you can remember that \((a+b)(a-b)=a^2-b^2\) always but you can also just multiply

OpenStudy (anonymous):

Oh! yes ok I remember now!

OpenStudy (misty1212):

the top is \[2(2a-9)(a+1)\] so lets multiply \[(a-9)(a+1)\] then by 2

OpenStudy (misty1212):

oops typo

OpenStudy (misty1212):

lets multiply \[(2a-9)(a+1)\] doing four multiplications, sometimes they say "foil"

OpenStudy (misty1212):

you want to try it first?

OpenStudy (anonymous):

yes I remember foil

OpenStudy (anonymous):

just a minute I'm doing it on paper, then I'll have you check it to see if its right, ok?

OpenStudy (misty1212):

kk

OpenStudy (anonymous):

wouldn't it be (2a-9) (2(a+1)) because of the 2 originally in front of a+1?

OpenStudy (misty1212):

yeah it would, but it would be easier to do \[(2a-9)(a+1)\] first and then double everything

OpenStudy (anonymous):

Oh does it come out the same?

OpenStudy (misty1212):

or you could just do \[(2a-9)(2a+2)\] if you prefer

OpenStudy (misty1212):

yes, multiplication is "associative" it doesn't matter which one you do first

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

4a2 - 14a - 18 ?

OpenStudy (anonymous):

4a^2 *

OpenStudy (misty1212):

yes!!

OpenStudy (misty1212):

final answer \[\frac{4a^2-14a-18}{a^2-9}\]

OpenStudy (anonymous):

yay! thank you so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]